OpenStudy (moazzam07):

A spinner is divided into many sections of equal size. Some sections are red, some are blue, and the remaining are green. The probability of the arrow landing on a section colored red is 12 over 20. The probability of the arrow landing on a section colored blue is 6 over 20. What is the probability of the arrow landing on a green-colored section? 2 over 20 6 over 20 8 over 20 18 over 20

OpenStudy (mathstudent55):

The probability of landing in a color is the number of sections of that color divided by all sections of all colors.

OpenStudy (moazzam07):

@mathstudent55 dont answer this i dont need the answer but thanks though

OpenStudy (mathstudent55):

\(P(red) = \dfrac{number ~of~red ~sections}{all~sections} = \dfrac{12}{20} \) \(P(blue) = \dfrac{number~of~blue ~sections}{all~sections} = \dfrac{6}{20} \) From the fractions above, what does it look like are the numbers of red sections, blue sections, and all sections?

OpenStudy (mathstudent55):

red = 12; blue = 6; all = 20 green = 20 - red - blue = 20 - 12 - 6 = 2 P(green) = 2/20